OpenStudy (anonymous):

.10. Explain how to solve 5x + 7 = 12 given the replacement set {0, 1, 2, 3, 4}.  Be clear in your explanation and show why your answer is correct. (Points : 4)

OpenStudy (anonymous):

replace x with each possible solution, in this case (0,1,2,3,4), find which ones equal 12.

OpenStudy (anonymous):

another medal......

OpenStudy (anonymous):

5(0) + 7 = 12 5(1) + 7 = 12 5(2) + 7 = 12 5(3) + 7 = 12 5(4) + 7 = 12

OpenStudy (anonymous):

sorry... it seems like im stocking you:)

OpenStudy (anonymous):

stalking*

OpenStudy (anonymous):

so @seiga what would be you're correct answer?

OpenStudy (anonymous):

im just waiting for an answer to my question in social science

OpenStudy (anonymous):

what would be you're solution set?

OpenStudy (anonymous):

oh, sorry ' stalking"

OpenStudy (anonymous):

@seiga ?

OpenStudy (anonymous):

5(0) + 7 = 7 5(1) + 7 = 12 5(2) + 7 = 17 5(3) + 7 = 22 5(4) + 7 = 27 Which one is true?

OpenStudy (anonymous):

I have a feeling @seiga is not going to answer lol

OpenStudy (anonymous):

i think seiga already got his answer

OpenStudy (anonymous):

i did.thanks though.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok cool